# Identification and Optimal Nonlinear Control of Turbojet Engine Using Koopman Eigenfunction Model

### v2

@@ 6.1.1 Data acquisition and processing
The training and test datasets were obtained from simulations using the in-house nonlinear CLM. Validation of this model against the commercial simulation software GasTurb was performed in [40]. The training dataset consisted of three 500-second segments. The first segment was formed by faster ramps with a rise time in range of 1-2 seconds. The command was changed approximately every 10 seconds. The second segment was a chirp signal with a frequency sweep from 0 to 0.25 Hz aiming to properly capture the low-frequency response. The last segment consisted of slow ramps with longer duration in the steady state, as it is crucial for the model to learn the steady-state regions well to avoid the prediction drifting away. The goal was to capture various aspects of the GTE dynamics. The test dataset was designed similarly but with overall faster dynamics. The sampling period was 10 ms for all simulations.
The smooth simulation output was further superimposed with Gaussian white noise with a standard deviation of 50 RPM generated using the MATLAB randn command. This way, the data characteristics got closer to those of real measurements. Furthermore, filtering of the datasets was performed using Savitzky-Golay polynomial filter. After subsequent normalization, the time derivatives of normalized spool speed were calculated using a 2nd-order finite-difference numerical differentiation scheme. The training and test datasets are depicted in Fig. 7(a) an Fig. 7(b), respectively.
@@ 6.1.2 SINDy
For identification of the GTE’s governing equation, the SINDy algorithm with logistic functions (LFs) was utilized. LFs are a popular choice for activation functions in neural networks and they have favorable properties [63]. As explained in Section 4.1, SINDy was combined with GD and ADAM optimizer and subjected to the L1 regularization of linear coefficients. Additionally, specifically for the LFs, and basically for any RBF, the sparsity could be further achieved also through the nonlinear parameters. Thus, these were optimized as well instead of just being guessed, or assumed to be evenly distributed, as is a common practice. The equation of LF is given as
@@
The following properties of LFs can help excluding some inactive or constant terms from the SINDy model:
@@
In the latter two cases, constants can be summed up to form a single constant function. The L1 regularization parameter was set to $10^{-4}$, and the zero derivative condition parameter was set to 0.005. The initial learning rates were set as 0.01, 50 and 0.01 for $\Xi$, $\varepsilon$ and $\mu$, respectively. The algorithm minimized a cost function given by the static $\dot{N}$ data. Additionally, every 10 iterations, the differential equation (23) was solved numerically for validation, and the true prediction MAE was stored. Once the prediction MAE started to grow, e.g. due to overfitting, the algorithm stopped.
The SINDy was initially performed for 5 LFs both for autonomous and input dynamics. Additional terms were given as $N,NW_{f},W_{f}$, yielding the total of 13 terms. After the first round of the GD, the first LF in $f(N)$, and the $NW_{f}$ term were excluded, and the rest of the terms were updated in the second round with smaller regularization parameter of $10^{-5}$. The resulting SINDy functions are shown in Fig. 8. It can be seen that the condition $f(0)=0$ was satisfied. Additionally, the steady-state fuel-flow characteristics were compared as shown in Fig. 9. For the SINDy model, the fuel flow was computed as $W_{f}=-f(N)/g(N)$ and de-normalized.
@@
Subsequently, the autonomous dynamics were extracted, and trajectories starting from multiple initial condition were generated as plotted in Fig. 10.
Figure 10: Autonomous trajectories of the GTE form the estimated $f(N)$ function
@@ 6.2 KEM construction from the SINDy model and the Kalman observer design
Based on the generated trajectories, the PSO was utilized to solve the constrained minimization problem (36) and to discover the optimal set of eigenvalues. Both the real and complex eigenvalues were considered and compared. The minimum and maximum real parts were set as -10 and -0.3, respectively, the maximum imaginary part was 10 and the maximum $\beta/\alpha$ ratio was 2.5. The PSO population size was 200 and the velocity weights were kept default.
Regarding the real eigenvalues, the automatic switching from distinct to repeated when necessary was enabled, and the repeated eigenvalues with secular terms led to better results. However, the complex eigenvalues outperformed the repeated eigenvalues and thus were utilized for the controller design. The system order was gradually increased from 4 to 8, and it was found that the 6th-order is an optimal trade-off between the complexity and accuracy.
@@
For each eigenfunction, a set of 20 LFs was optimized. However, it was later found out that just 10 LFs were also sufficient. This reveals yet another advantage over DMD-based approaches – as the basis functions can adapt to each eigenfunction individually, eigenfunctions can have different number of terms, depending on the complexity of the eigenfunction’s shape. Furthermore, the basis can be decided based on the discovered shapes avoiding the trial-and-error selection for EDMD.
The resulting eigenfunctions are shown in Fig. 12(a). Subsequently, the vector of mode amplitudes was found using Eq. (51) (or Eq. (52)). The comparison of repeated real and complex eigenvalues is depicted in Fig. 12(b).
Figure 12: (a) The resulting eigenfunctions, and (b) the comparison of projection results for the repeated real and complex eigenvalues
[…]
@@
Despite a different order, in which the control system design was presented in Section 5.4, here the Kalman observer design is described. The covariance matrix of the eigenfunction prediction error was computed and is depicted in Fig. 14(a). It can be seen that there are off-diagonal terms that are not negligible. This matrix was utilized for tuning of the process noise weighting matrix $\mathbf{Q_{o}}$. The measurement noise covariance R was estimated from a steady-state sample taken from unfiltered normalized spool-speed data and its value is also shown in the figure. The performance of KEM with the Kalman observer with MAE and MAPE is depicted in Fig. 14(b). Both the metrics were calculated based on the difference between the estimation and clean data (with no noise). It can be concluded that the Kalman observer provides very accurate estimation of both the eigenfunctions (directly) and also the spool speed (indirectly).
@@
Figure 15 depicts the residual histograms. These are evaluated for a) the prediction using the baseline KEM model for the filtered test dataset (red), b) the prediction of the noisy test dataset using the Kalman observer, which reveals the Gaussian distribution of the added white noise (green), and c) residuals given by the difference between the prediction of the Kalman observer and the clean test dataset (with no added noise) to assess the estimation performance (blue).
Figure 15: Residual histograms of baseline KEM without KF correction for filtered test dataset, noisy prediction using KF that reveals the Gaussian white noise distribution with variance of 50 RPM, and estimation error computed using clean data without noise
@@ 6.2.1 Thrust estimation
As also the direct thrust control (DTC) belongs to modern research topics of high interest, the normalized corrected thrust from another test dataset was projected onto the span of $\hat{\beta}(t)$. The thrust output equation and the corresponding LS problem are given as
@@
$\mathbf{C_{F}}=\arg$ $\min_{\mathbf{C_{F}}}\|F_{c,n}-\mathbf{C_{F}}\mathbf{\Phi}-\mathbf{D_{F}}W_{f}\|_{2}^{2},$ (97)
where $\mathbf{C_{F}}$ is a vector of projection coefficients, and $F_{c,n}$ is the normalized corrected thrust given as
@@
The thrust prediction results are shown in Fig. 16. The MAPE is inherently higher because the lower thrust boundary is near 0 and thus the relative error grows exponentially for the same MAE. The results demonstrate that the KEM can also be effectively utilized for DTC. In this paper however, only the control of spool speed is considered.
@@ 6.3.1 Classical and gain scheduled nonlinear PI controllers
[…]
For the LPV-PI design, the discovered SINDy model was analytically differentiated, and the engine gain and time constant were assessed for a set of 11 evenly distributed steady-state points. For each steady-state point, the PI controller was initially designed using the MATLAB pidtune function and further fine-tuned manually in the Live Editor app Tune PID Controller to reach the demanded performance with minimum overshoots. The PI gains were fitted by 5th-order polynomials as depicted in Fig. 17.
@@ 6.3.2 SINDy - IMC
The filter time constant was manually tuned to $\tau_{f}=0.05$ s. Lower values significantly increased the sensitivity to the measurement noise, whereas higher values worsened the tracking performance. The resulting controller gains of the PID part are plotted in Fig. 18. It can be observed that the trends of proportional and integral gains are similar to that of the LPV-PI controller.
@@ 6.3.3 Koopman LQGI controller
The LQGI controller was designed using the proposed approach with GA or PSO from the MATLAB Global Optimization Toolbox. GA yielded better results after multiple initializations. Regarding the hyperparameters, the population size and number of generations were both set to 100, the elite count was set to 3, and the replication-crossover-mutation rates and parameters were kept default. The initial population was generated using a Latin hypercube sampling function. The resulting optimal $\mathbf{Q_{\varphi}}$ matrix is plotted in Fig. 19(a). The $\mathbf{Q_{f}}$ and $\mathbf{Q_{i}}$ values are also shown, and it can be seen that the algorithm generated a rather aggressive integral gain possibly balanced by the $\mathbf{dQ_{\varphi}}$ matrix targeting individual modes. The gain and phase margins satisfied the limits as shown in Fig. 19(b). Additionally, the disk-based margins were evaluated using MATLAB Robust Control Toolbox, and the minimum disk-based GM and PM were $\pm$9.33 dB and $\pm$52.4 deg, respectively. Variation of closed-loop eigenvalues with spool speed in Fig. 20(a) points at the closed-loop system being stable. The Bode plot of sensitivity, complementary sensitivity, and open-loop transfer functions in Fig. 21 shows that the system also possessed high reference tracking capability and robustness in the typical engine control bandwidth of 10 rad/s [64]. This was further validated by disturbance rejection testing. The peak sensitivity was 3.22 dB and was located outside of this bandwidth.
@@
Figure 20: (a) The range of eigenvalues of the closed-loop system with changing spool speed and (b) the $\mathbf{K_{\beta}}$ gains
@@ 6.4.1 Reference tracking
The performance of controllers was evaluated using the nonlinear CLM at sea-level and varying flight conditions. The command profile was the same stair profile used for the optimization of K-LQGI in Fig. 4. The results for sea-level conditions are depicted in Fig. 22 and details in Fig. 23. The wIAE in the figures stands for the weighted IAE in Eq. (71). Since the LPV-PI and IMC controllers were based on only an approximation of the engine nonlinear dynamics valid in the vicinity of the steady-state operating line, the different dynamical characteristics of acceleration and deceleration were handled poorly. By further increase of LPV controller gains, the response was be faster also during deceleration, but the overshoots became unacceptable during acceleration and small transients. Similarly, further decreasing of the filter time constant $\tau_{f}$ of the IMC resulted in faster response at the cost of significantly increased sensitivity of the derivative term to the measurement noise.
Quantitatively, the K-LQGI controller with the Kalman observer exhibited the lowest wIAE. The performance during small transients is also of high significance, and the K-LQGI kept the same response characteristics, namely the speed and minimal overshoots, across the entire operating range, whereas the LPV-PI controller presented an overall slower response. The IMC controller got closest to the K-LQGI performance but was also overall slower. Regarding the large transients, the initial phase was dominantly guided by the fuel flow limiters. Thus, the performance of LPV-PI, IMC and K-LQGI controllers was very similar, except the deceleration to GI, where both LPV-PI and IMC controllers exhibited a sudden decrease of the response speed with decrease of the spool speed. This points at the nonlinear nature of the engine dynamics, further supported by the fuel flow limiting, which could not be captured by the scheduling approach. Since limiters were inactive during small transients, the fact that K-LQGI did not cause overshoots points at the ability of the controller to decrease the acceleration rate close to the desired state accordingly. The Kalman estimates showed the ability of KEM to predict the engine state reliably.
@@
Figure 23: Details of certain response regions - (a) Small transient for low RPM acceleration, (b) high RPM, (c) low RPM deceleration, (d) large transient - acceleration and (e) deceleration
Concerning the varying flight conditions, the altitude and MN profiles are plotted in Fig. 24. These changes are unrealistic, but the main goal was to capture the effects of flight conditions during the 70-second simulation period. The results are depicted in Fig. 25. The classical and LPV PI controllers started to lose robustness not being able to effectively reject the corresponding disturbances in engine behavior. The IMC performance was consistent but slightly slower compared to the K-LQGI. The deceleration to GI was not considered for the comparison during the higher MN of flight, as the minimum sustainable spool speed increases with MN, and the simulation would leave the known region of the compressor map leading to nonphysical simulation outputs.
@@
The wIAE, standard IAE and maximum settling time (MST) are summarized in Table 1. The MST was defined as a time when the response last crossed the $N_{d}\pm 1\%$ boundary during the full acceleration. The quantitative results also point at the superior performance of the K-LQGI controller closely followed by the SINDy-IMC controller.

### v6

@@ 6.1.1 Data acquisition and processing
The training and test datasets were obtained from simulations using the in-house nonlinear CLM. Validation of this model against the commercial simulation software GasTurb was performed in [bib40]. The training dataset consisted of three 500-second segments. The first segment was formed by faster ramps with a rise time in the range of 1-2 seconds. The command was changed approximately every 10 seconds. The second segment was a chirp signal with a frequency sweep from 0 to 0.25 Hz, aiming to properly capture the low-frequency response. The last segment consisted of slow ramps with longer duration in the steady state, as the model must learn the steady-state regions well to avoid the prediction drifting away. The goal was to capture various aspects of the GTE dynamics. The test dataset was designed similarly, but with overall faster dynamics. The sampling period was 10 ms for all simulations.
To simulate the measured data, the smooth simulation output was further superimposed with Gaussian white noise with a standard deviation of 50 RPM generated using the MATLAB randn command. This slightly exceeds the typical RPM-sensor uncertainty standard deviation reported to be $0.25$ $\%$ of nominal RPM for gas turbines [Wei2025]. Furthermore, filtering of the datasets was performed using a Savitzky-Golay polynomial filter. After subsequent normalization, the time derivatives of normalized spool speed were calculated using a 2nd-order finite-difference numerical differentiation scheme. The training and test datasets are depicted in Fig. 8(a) and Fig. 8(b), respectively.
@@ 6.1.2 SINDy
For identification of the GTE’s governing equation, the SINDy algorithm with logistic functions (LFs) was utilized. LFs are a popular choice for activation functions in neural networks, and they have favorable properties [bib63]. As explained in Section 4.1, SINDy was combined with GD and ADAM optimizer and subjected to the L1 regularization of linear coefficients. Additionally, specifically for the LFs and basically for any RBF, the sparsity could be further achieved through the nonlinear parameters. Thus, these were optimized as well instead of just being guessed, or assumed to be evenly distributed, as is a common practice. The equation of LF is given as
@@
The following properties of LFs can help exclude some inactive or constant terms from the SINDy model:
@@
In the latter two cases, constants can be summed up to form a single constant function. The L1 regularization parameter was set to $10^{-4}$, and the zero derivative condition parameter was set to 0.005. The initial learning rates were set as 0.01, 50, and 0.01 for $\Xi$, $\varepsilon$, and $\mu$, respectively. The algorithm minimized a cost function given by the static $\dot{N}$ data. Additionally, every 10 iterations, the differential equation (25) was solved numerically for validation, and the true prediction MAE was stored. Once the prediction MAE started to grow, e.g., due to overfitting, the algorithm stopped.
The SINDy was initially performed for 5 LFs, both for autonomous and input dynamics. Additional terms were given as $N,NW_{f},W_{f}$, yielding a total of 13 terms. After the first round of the GD, the first LF in $f(N)$, and the $NW_{f}$ term were excluded, and the rest of the terms were updated in the second round with a smaller regularization parameter of $10^{-5}$. The resulting SINDy functions are shown in Fig. 9. It can be seen that the condition $f(0)=0$ was satisfied. Additionally, the steady-state fuel-flow characteristics were compared as shown in Fig. 10. For the SINDy model, the fuel flow was computed as $W_{f}=-f(N)/g(N)$ and de-normalized.
@@
Subsequently, the autonomous dynamics were extracted, and trajectories starting from multiple initial conditions were generated as plotted in Fig. 11.
Figure 11: Autonomous trajectories of the GTE from the estimated $f(N)$ function
@@ 6.2 KEM construction from the SINDy model and the Kalman observer design
Based on the generated trajectories, the PSO was utilized to solve the constrained minimization problem (38) and to discover the optimal set of eigenvalues. Both the real and complex eigenvalues were considered and compared. The minimum and maximum real parts were set as -10 and -0.3, respectively; the maximum imaginary part was 10, and the maximum $\beta/\alpha$ ratio was 2.5. The PSO population size was 200, and the velocity weights were kept default.
Regarding the real eigenvalues, the automatic switching from distinct to repeated when necessary was enabled, and the repeated eigenvalues with secular terms led to mean absolute error (MAE) of $0.001$. However, the complex eigenvalues outperformed the repeated eigenvalues with MAE of $4.5\times 10^{-4}$, and thus were utilized for the controller design. The system order was gradually increased from 4 to 8, and it was found that the 6th-order is an optimal trade-off between the complexity and accuracy.
@@
Originally, a set of 20 LFs was optimized for each eigenfunction. However, it was later found out that just 10 LFs were also sufficient. This reveals yet another advantage over DMD-based approaches – as the basis functions can adapt to each eigenfunction individually, eigenfunctions can have a different number of terms, depending on the complexity of the eigenfunction’s shape. Furthermore, the basis can be decided based on the discovered shapes, avoiding the trial-and-error selection for EDMD.
The resulting eigenfunctions are shown in Fig. 13. Subsequently, the vector of mode amplitudes was found using (54) (or (55)).
Figure 13: The resulting eigenfunctions
Following the identification, the final model was validated using the test dataset. The validation was performed for the filtered data to assess the “off-line” prediction accuracy, which is not affected by noise. The MAE and the mean absolute percentage error (MAPE) were utilized as the error metrics. They were computed for the de-normalized corrected spool speed. For the training data, the final MAE and MAPE were 28.44 RPM and 0.33 $\%$, respectively. The test dataset prediction results, MAE, and MAPE are depicted in Fig. 14.
@@
Subsequently, the Kalman observer was designed. The covariance matrix of the eigenfunction prediction error was computed and is depicted in Fig. 15(a). It can be seen that there are off-diagonal terms that are not negligible. This matrix was utilized for tuning of the process noise weighting matrix $\mathbf{Q_{o}}$. The measurement noise covariance $\mathbf{R_{o}}$ was estimated from a steady-state sample taken from unfiltered normalized spool-speed data, and its value is also shown in the figure. The performance of KEM with the Kalman observer with MAE and MAPE is depicted in Fig. 15(b). Both metrics were calculated based on the difference between the estimation and clean data (with no noise). It can be concluded that the Kalman observer provides very accurate estimation of both the eigenfunctions (directly) and also the spool speed (indirectly).
@@
Figure 16 depicts the residual histograms. These are evaluated for a) the prediction using the baseline KEM model for the test dataset filtered using the Savitzky-Golay filter (red), b) the prediction of the noisy test dataset using the Kalman observer, which reveals the Gaussian distribution of the added white noise (green), and c) residuals given by the difference between the estimation of the Kalman observer and the clean test dataset (with no added noise) to assess the estimation performance (blue).
Figure 16: Residual histograms of the baseline KEM without the KF for the dataset filtered using the Savitzky-Golay filter (red), error of prediction of the noisy dataset using the KF that reveals the Gaussian white noise distribution with variance of 50 RPM (green), and estimation error computed using clean data without noise (blue)
@@ 6.2.1 Thrust estimation
As the direct thrust control also belongs to modern research topics of high interest, the normalized corrected thrust from another test dataset was projected onto the span of $\hat{\boldsymbol{\zeta}}(t)$. The thrust output equation and the corresponding LS problem are given as
@@
$\begin{bmatrix}\mathbf{C_{F}}&\mathbf{D_{F}}\end{bmatrix}=\arg$ $\min_{\mathbf{C_{F},D_{F}}}\|F_{c,n}-\mathbf{C_{F}}\mathbf{\Phi}-\mathbf{D_{F}}W_{f}\|_{2}^{2},$ (101)
where $\mathbf{C_{F}}$ and $\mathbf{D_{F}}$ are vectors of coefficients of projection onto $\mathbf{\hat{\Phi}}$ and $W_{f}$, respectively, and $F_{c,n}$ is the normalized corrected thrust given as
@@
The thrust prediction results are shown in Fig. 17. The MAPE is inherently higher because the lower thrust boundary is near 0, and thus the relative error grows exponentially for the same MAE. The results demonstrate that the KEM can also be effectively utilized for thrust control. In this paper, however, only the control of the spool speed is considered.
@@ 6.3.1 Classical and gain scheduled nonlinear PI controllers
Next, the controllers were designed. The classical PI controller was tuned to satisfy the aforementioned requirements arising from the PART 31 regulation, stating that the engine should be able to respond to a 1-second linear ramp command from 15 $\%$ to 95 $\%$ of the maximum thrust in under 5 seconds. The resulting gains were given as $K_{p}=1$ and $K_{i}=3$.
For the LPV-PI design, the discovered SINDy model was analytically differentiated, and the engine gain and time constant were assessed for a set of 11 evenly distributed steady-state points. At each steady-state point, the PI controller was tuned to reach a fast step response with no overshoot, minimizing the integral of quadratic tracking error. To avoid unreasonable gain values, additional regularization of the gains was included. The optimal PI gains were fitted by 5th-order polynomials as depicted in Fig. 18.
@@ 6.3.2 SINDy - IMC
The filter time constant was manually tuned to $\tau_{f}=0.05$ s. Lower values significantly increased the sensitivity to the measurement noise, whereas higher values worsened the tracking performance. The resulting controller gains of the PID part are plotted in Fig. 19. It can be observed that the trends of proportional and integral gains are similar to those of the LPV-PI controller.
@@ 6.3.3 Koopman LQGI controller
The LQGI controller was designed using the proposed approach with GA or PSO from the MATLAB Global Optimization Toolbox. GA yielded better results after multiple initializations. Regarding the hyperparameters, the population size and number of generations were both set to 100, the elite count was set to 3, and the replication-crossover-mutation rates and parameters were kept at the default. The initial population was generated using a Latin hypercube sampling function. The resulting optimal $\mathbf{Q_{\Phi}}$ matrix is plotted in Fig. 20(a). The $\mathbf{Q_{f}}$ and $\mathbf{Q_{i}}$ values are also shown, and it can be seen that the algorithm generated a rather aggressive integral gain, possibly balanced by the $\mathbf{dQ_{\Phi}}$ matrix targeting individual modes. The gain and phase margins satisfied the limits, (83) and (84), as shown in Fig. 20(b). Additionally, the disk-based margins were evaluated using MATLAB Robust Control Toolbox, and the minimum disk-based GM and PM were $\pm$9.33 dB and $\pm$52.4 deg, respectively. Variation of closed-loop eigenvalues with spool speed in Fig. 21(a) points to the closed-loop system being stable. The Bode plot of sensitivity, complementary sensitivity, and open-loop transfer functions in Fig. 22 shows that the system also possessed high reference tracking capability and robustness in the typical engine control bandwidth of 10 rad/s [bib64]. This was further validated by disturbance rejection testing. The peak sensitivity was 3.22 dB and was located outside of this bandwidth.
@@
Figure 21: (a) The range of eigenvalues of the closed-loop system with changing spool speed and (b) the $\mathbf{K_{\boldsymbol{\zeta}}}$ gains
@@ 6.4.1 Reference tracking
The performance of controllers was evaluated using the nonlinear CLM at sea level and varying flight conditions. The command profile was the same stair profile used for the optimization of K-LQGI in Fig. 5. The results for sea-level conditions are depicted in Fig. 23 and details in Fig. 24. The wIAE in the figures stands for the weighted IAE in (5.4.2). Since the LPV-PI and IMC controllers were based on only an approximation of the engine’s nonlinear dynamics valid in the vicinity of the steady-state operating line, the different dynamical characteristics of acceleration and deceleration were handled poorly. By further increasing the LPV controller gains, the response was faster also during deceleration, but the overshoots became unacceptable during acceleration and small transients. Similarly, further decreasing the filter time constant $\tau_{f}$ of the IMC resulted in faster response at the cost of significantly increased sensitivity of the derivative term to the measurement noise.
Quantitatively, the K-LQGI controller with the Kalman observer exhibited the lowest wIAE. The performance during small transients is also of high significance, and the K-LQGI kept the same response characteristics, namely the speed and minimal overshoots, across the entire operating range, whereas the LPV-PI controller presented an overall slower response. The IMC controller got closest to the K-LQGI performance, but was also overall slower. Regarding the large transients, the initial phase was dominantly guided by the fuel flow limiters. Thus, the performance of the LPV-PI, IMC, and K-LQGI controllers was very similar, except for the deceleration to GI, where both LPV-PI and IMC controllers exhibited a sudden decrease in response speed, accompanied by a decrease in spool speed. This points to the nonlinear nature of the engine dynamics, further supported by the fuel flow limiting, which could not be captured by the scheduling approach. Since limiters were inactive during small transients, the fact that K-LQGI did not cause overshoots indicates the controller’s ability to decrease the acceleration rate close to the desired state accordingly. The Kalman estimates showed the ability of KEM to predict the engine state reliably.
@@
Figure 24: Details of certain response regions - (a) Small transient for low RPM acceleration, (b) high RPM, (c) low RPM deceleration, (d) large transient - acceleration. Solid lines were computed using the CLM without noise. The black dashed line is the estimation for noisy data using the Kalman observer
Concerning the varying flight conditions, the altitude, $H$, and $M_{0}$ profiles are plotted in Fig. 25. These changes are unrealistic, but the main goal was to capture the effects of flight conditions during the 70-second simulation period. The results are depicted in Fig. 26. The classical and LPV PI controllers started to lose robustness, not being able to effectively reject the corresponding disturbances in engine behavior. The IMC performance was consistent but slightly slower compared to the K-LQGI. The deceleration to GI was not considered for the comparison during the higher $M_{0}$ of flight, as the minimum sustainable spool speed increases with $M_{0}$, and the simulation would leave the known region of the compressor map, leading to nonphysical simulation outputs.
@@
The wIAE, standard IAE, and maximum settling time (MST) are summarized in Table 1. The MST was defined as a time when the response last crossed the $N_{d}\pm 1\%$ boundary during the full acceleration. The quantitative results also point to the superior performance of the K-LQGI controller, closely followed by the SINDy-IMC controller.
@@
6.4.2 Disturbance rejection
The most common disturbances acting on the engine are caused by the changing flight conditions - $H$ and $M_{0}$, and aircraft maneuvers affecting the air flow conditions, e.g., increased transient flow separation in the intake. Here, the disturbance rejection capability of the compared controllers was assessed for unrealistic abrupt changes in flight conditions, depicted in Fig. 27, leading to changes in the inlet total pressure and temperature. The setpoint was set to 90 $\%$ of the maximum spool speed as it often corresponds to the nominal operating point, where the engine spends most of the time. The results plotted in Fig. 28 show that the K-LQGI controller exhibited the best disturbance rejection, with the SINDy-IMC close by.
Figure 27: Disturbances modeled by aggressive changes in (a) altitude and (b) Mach number
Figure 28: Disturbance rejection results. The K-LQGI controller exhibits the best disturbance rejection (red), followed by the SINDy-IMC (magenta)
7 Conclusion
To the author’s best knowledge, based on a rather extensive literature review, this is the first paper utilizing the KOT for the design of an optimal GTE controller. Furthermore, it is also the first to combine the SINDy algorithm with the KOT for estimation of the GTE governing equation with a subsequent Koopman canonical transform to the KEM. Another contribution is the constrained optimization of a low-order set of eigenvalues using metaheuristic global optimization tools, e.g., GA and PSO, and the projection of the autonomous trajectories in the time domain, providing an optimal set of eigenvalues for the eigenfunction construction from data. The objective function rejects the unstable, too oscillatory, and too lightly damped eigenvalues. Moreover, the algorithm can switch from distinct eigenvalues to repeated ones when the distance of any two eigenvalues is smaller than a specified margin. The discovered set of eigenfunctions and eigenvalues can also be extended by exploiting the algebraic structure. However, by a product of two eigenfunctions, we obtain a new eigenfunction that always decays faster than the original two functions, and thus it contributes less to the dynamics.
The KEM with the Kalman observer was capable of an accurate state prediction even outside the region of validity of the LPV system. Furthermore, it allows for more aggressive integral control balanced by targeting of individual modes, where the modes with the lowest damping are the most important ones. This is captured by the reduced overshoots compared to PI and LPV-PI controllers. Furthermore, the LPV-PI and IMC controllers both struggle with the large deceleration, which points to too small PI gains in the low-RPM region for the deceleration. However, further increases of these gains would cause larger overshoots during small accelerations for the LPV-PI controller and increased sensitivity to noise for the IMC controller. Thus, the locally linearized controllers are overall effectively outperformed by the optimized KEM LQGI.
These methods can be extended to an adaptive framework. For the KEM, the eigenfunctions can be fixed, and the matrix of Koopman modes, $\mathbf{C}$, can be adapted in real time using the ridge regression for a finite sequence of past spool speed measurements. Also, the control law can be further replaced, e.g., by a nonlinear MPC controller. Both the identification method and the subsequently designed controller proved to be efficient for the control of a single-spool turbojet GTE. For multi-spool turbojet and turbofan GTEs, the proposed eigenvalue optimization needs to be modified. This is currently the subject of our ongoing research. The issue is primarily the ambiguity in coefficients K encoding both the unknown initial conditions of eigenfunctions and the mode amplitudes. These cannot be easily decoupled like in the case of 1st-order systems, where the variation of K for different trajectories is caused solely by the changing initial conditions of $\mathbf{\Phi}$. Fortunately, the presented eigenfunction identification procedure works in higher dimensions as well.
In future research, we will extend this framework to multi-rotor and turbofan engines and test other control strategies such as the MPC.